Suppose,demand function for and supply function of brown price are Qd=30-2P and Qs=20+3P respectiovely.
1)Calculate the equilibrium price and quantity of brown rice.
2)Suppose,due to the campaign that consumption of brown rice lowers cholesterol level in human body,the demand for brown rice has been increased to Qd=50-2P.What kind of impact of the campaign will be on the market price of brown rice?what will happen to the market price of plain rice?
1
Expert's answer
2014-05-23T12:18:18-0400
Qd=30-2P and Qs=20+3P respectively.

1)Calculate the equilibrium price and quantity of brown rice.
In equilibrium Qd = Qs, so:
30 - 2P = 20 + 3P
5P = 10
Pe = $2
Qe = 30 - 2*2 = 26 units.

2)If the demand for brown rice has been increased to Qd=50-2P, the market price
of brown rice will increase to. To find new equilibrium price we should equal
new Qd = Qs.
50 - 2P = 20 + 3P
5P = 30
Pe = $6.
Qe = 50 - 2*6 = 38 units.